# Investigating the Impact of Trust in Multi-Human Multi-Robot Task Allocation

### v1

@@ III-B Simulation
Humans: $v_{H}=1.0$ m/s, $s_{H}=10$ m; UAVs: $v_{UAV}=2.0$ m/s, $s_{UAV}=20$ m; UGVs: $v_{UGV}=1.5$ m/s, $s_{UGV}=15$ m. These speeds are chosen to reflect realistic movement capabilities in SAR. The simulation progresses in discrete time steps, where each step represents 1 second of real-time. We implemented and compared five trust models: 1) NoTrust: A baseline model with no trust updates, 2) Monir: Based on Monir et al.’s work, using discrete trust regions, 3) XuDudek: Implementing Xu and Dudek’s continuous trust update mechanism, 4) GuoYang: Utilizing Guo and Yang’s Beta distribution-based trust model, and 5) ECT: Our proposed ECT model incorporates task complexity and robot capability. Tasks are allocated using a Q-learning approach [43]. The state space $S$ is defined by the agent’s distance to the nearest POI and its current trust level. The action space $A$ consists of three possible actions: move to POI, perform task, and wait. The Q-learning update rule is:
$Q(s,a)\leftarrow Q(s,a)+\alpha[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)]$ (10)
where $\alpha=0.1$ is the learning rate, $\gamma=0.9$ is the discount factor, and $r$ is the reward. The exploration rate $\epsilon$ is set to 0.1.
We conducted a series of experiments to evaluate the performance of different trust models across various team compositions and environmental conditions, including A) Small team (2 Humans and 2 Robots; 2H-2R) in flat terrain, B) Medium team (5 Humans and 5 Robots; 5H-5R) in rough terrain, and C) Large team (10 Humans and 10 Robots; 10H-10R) in obstacle-dense terrain. The environments types (e.g., flat terrain) represent levels of task complexity. We used two primary metrics to evaluate performance, including:
@@
    $TSR=\frac{\text{Number of completed tasks}}{\text{Total number of tasks}}$ (11)
@@
    $ACT=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (12)
where $t_{i}$ is the completion time of task $i$, and $n$ is the total number of completed tasks.
For each combination of trust model and scenario, we conducted 100 independent simulation runs to ensure statistical significance. Each run lasted for a maximum of 500 time steps or until all tasks were completed. We used the trust questionnaire from Yagoda and Gillian [44] as our theoretical framework for defining trust parameters, specifically focusing on measuring function and performance-related trust parameters, which include competence, ability, and dependability. The simulation logged key metrics at each step, including agent positions, trust levels, task completion times, and task success rates. This data allowed for a comparison of the trust models in terms of their impact on team performance and collaboration in a multi-agent SAR environment.
@@ IV Results and Analysis
(a) Success rate
(b) Average of completion time
Figure 4: Results of the Success Rate and Average Completion Time of the Different Trust Models in team compositions. The star symbol indicates the best performance in each result, the highest success rate, and the shortest completion time. Zoom in for better readability.
Results from our evaluation of the five trust models, including the No Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, and the ECT Model across three team compositions (e.g., 2H-2R, 5H-5R, and 10H-10R), revealed close results in both task success rates and average task completion times. In small team configurations (2H-2R), we observed that all trust models performed comparably in terms of success rates. ECT, NoTrust, and XuDudek all achieved the highest success rate of 0.35 ± 0.08, while Monir and GuoYang followed closely at 0.33 ± 0.08. Regarding completion times, XuDudek was the fastest at 49.28 ± 16.72 seconds, with ECT following closely at 49.66 ± 14.43 seconds. Notably, ECT demonstrated the lowest standard deviation in completion time, indicating more consistent performance even in small-scale operations. As we moved to medium-sized teams (5H-5R), the performance differences between trust models became more pronounced. Monir took the lead in success rate at 0.63 ± 0.10, with ECT following closely at 0.62 ± 0.10. The other models (NoTrust, XuDudek, and GuoYang) tied at 0.60, albeit with slightly higher standard deviations. Interestingly, the completion time results for this team size revealed a different pattern. Monir was the fastest at 48.42 ± 11.73 seconds, while ECT had the slowest mean completion time at 52.25 ± 11.25 seconds. This suggests a potential trade-off in ECT between maintaining a high success rate and optimizing completion time in moderately complex scenarios.
The most notable results emerged in the large team configuration (10H-10R), where ECT demonstrated its strengths. In this complex scenario, ECT tied with GuoYang for the highest success rate at 0.82 ± 0.10, outperforming the other models. Moreover, ECT achieved the fastest mean completion time of 48.99 ± 9.13 seconds, significantly quicker than the next fastest model, Monir, at 49.14 ± 10.08 seconds. Also, the ECT model exhibited the lowest standard deviations in both success rate and completion time, indicating the most consistent and reliable performance in large-scale MH-MR contexts. An important observation from these results is the scalability demonstrated by the ECT model as its success rate improved from 0.35 in small teams to 0.82 in large teams, representing one of the most significant improvements among all models. This scalability suggests that the ECT model’s trust dynamics become effective as team complexity grows, a critical factor for real-world applications where scalability is often paramount. In the 10H-10R configuration, the ECT model outperformed other models in terms of consistency, demonstrating the lowest standard deviations in both success rate (±0.10) and completion time (±9.13s).
V Discussion
Non-Linearity in Team Size and Performance. Results from our experiment showed a non-linear relationship between team size and performance across the different trust models. This non-linearity aligns with findings from prior work [45] which showed that large autonomous systems often exhibit non-linear behaviors due to the complex interactions between individual robots. The superior performance of trust-based models, particularly the ECT model, in larger team configurations (10H-10R) suggests a potential ’scale threshold’ where trust becomes a critical factor in team efficiency and optimal performance. The findings from our research open up new avenues for research into the optimal team sizing and threshold where trust measurement becomes crucial for MH-MR collaborations.
Trust as a Computational Resource. The ECT’s model’s performance, especially in larger teams introduces a novel perspective of trust as a form of distributed computational resource that leverages the perception of team members vis-a-vis robot performance. Unlike traditional views of trust as a purely psychological construct, our model and the results from the experiment suggest that trust functions as a form of collective intelligence that scales with team size and requires active participation from all team members for an objective outcome. This conceptualization of ’trust as a form of computation’ could help researchers reimagine how we design and implement trust models in robotic systems, potentially leading to an easily scalable and adaptable approach for eliciting and computing trust for large MH-MR teams. The design of the ECT model implicitly means that as the team size increases, the computational power of trust also increases.
Task Allocation and Trust Heuristics in Large Teams. The performance discrepancies between trust-based and non-trust models were more pronounced in larger team configurations and raised important questions about the role of trust in mitigating complexity in MH-MR teaming and task allocation. As team size increases, the computational demands of task allocation grow as well. Our results suggest that trust might serve as an additional heuristic, in addition to other factors like fatigue and workload, to reduce this complexity by providing insight and allowing for more efficient task allocation that is based on empirical metrics in large-scale MH-MR collaboration settings. Specifically, the ECT model is a scalable trust model that is adaptive to different team configurations and sizes and performs relatively well regardless of team composition, showing its usefulness.
Advantages of the ECT Model. Another key benefit of the ECT model lies in its direct approach. While other trust models often rely on independent techniques like facial detection, behavioral cues, and other indirect and after the fact approaches, the ECT model relies on direct elicitation of trust scores from human team members for a task and continuously throughout the lifecycle of the task. This approach ensures that the trust levels of human teams on robot performance are continuously updated. Furthermore, the approach of team trust aggregation allows for congregating the trust scores of all team members in measuring trust levels. Through this approach, the trust rating of an individual team does not have an outsized influence on the overall trust final score generated by the team.
Study Limitations and Future Directions. While results from our study provided valuable insights on trust-based task allocation in MH-MR teams, it is important to also acknowledge the limitations of our study. First, we used simulated environments for our experiment. While this allows for controlled experiments that do not expose users to harm, it may not fully capture the complexities that are often present in real-world contexts. Furthermore, our task domain might not translate to other scenarios and domains. As per prior work, trust is not easy to model, and as a result, our direct question-based method might not capture all the relevant nuances required for measuring trust. Despite these limitations, our work provides a good foundation for future research in trust-based task allocation approaches for MH-MR teams.
VI Conclusion
In this paper, we investigated the impact of trust on task allocation in multi-human multi-robot teams. We employed five trust models, including N-Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, and the ECT Model across different types of teams that included 2H-2R, 5H-5R, and 10H-10R. Our simulation task involved an SAR activity. Findings from our research revealed that trust-based models outperformed the no-trust model most of the time in the 10H10R team configuration and that there was no noticeable difference in the lower configurations. These findings have implications for trust in the context of MH-MR task allocation. Future work could explore trust repair mechanisms for MH-MR teams where trust has been broken due to poor task performance.

### v2

@@ III-B Simulation
Humans: $v_{H}=1.0$ m/s, $s_{H}=10$ m; UAVs: $v_{UAV}=2.0$ m/s, $s_{UAV}=20$ m; UGVs: $v_{UGV}=1.5$ m/s, $s_{UGV}=15$ m. These speeds are chosen to reflect realistic movement capabilities in SAR. The simulation progresses in discrete time steps, where each step represents 1 second of real-time. We implemented and compared seven trust models: 1) NoTrust: A baseline model with no trust updates, 2) Monir: Based on Monir et al.’s work, using discrete trust regions, 3) XuDudek: Implementing Xu and Dudek’s continuous trust update mechanism, 4) GuoYang: Utilizing Guo and Yang’s Beta distribution-based trust model, 5) Ali et al. Bounds-Based Trust Model, Guo et al. Beta-Distribution Trust Model, 7) ECT: Our proposed ECT model incorporates task complexity and robot capability. Tasks are allocated and executed via a Q-learning approach [33], where each agent chooses among three actions: move to the POI, perform the task, or wait. We define the state space
$S=\bigl{(}\text{distance to nearest POI},\,\text{current trust level},\\ \text{POI complexity},\,\text{skill match}\bigr{)},$ (19)
and the action space
$A=\{\text{move to POI},\,\text{perform task},\,\text{wait}\}.$
At each step, the agent observes the current state $s\in S$, takes an action $a\in A$, and receives a reward $r$ (e.g., positive for successful task completion, negative for failures or wasted movement). We use the standard Q-learning update rule:
$Q(s,a)\;\leftarrow\;Q(s,a)\;+\;\alpha\Bigl{[}r+\gamma\,\max_{a^{\prime}}\,Q(s^{\prime},a^{\prime})\;-\;Q(s,a)\Bigr{]},$ (20)
where $\alpha=0.1$ is the learning rate, $\gamma=0.9$ is the discount factor, and $s^{\prime}$ is the next state after executing action $a$. We employ an $\epsilon$-greedy policy with $\epsilon=0.1$ for exploration, meaning agents choose a random action with probability $0.1$ and the Q-learning action otherwise.
Figure 3: Illustration of our virtual simulation environment containing different agents (UGVs, UAVs, and humans) and different points of interests.
We conducted a series of experiments to evaluate the performance of different trust models across various team compositions and environmental conditions, including A) Small team (2 Humans and 2 Robots; 2H-2R) in flat terrain, B) Medium team (5 Humans and 5 Robots; 5H-5R) in rough terrain, and C) Large team (10 Humans and 10 Robots; 10H-10R) in obstacle-dense terrain. The environment types (e.g., flat terrain) represent levels of task complexity. We used two primary metrics to evaluate performance, including:
@@
  • Agent Utilization: Tracks the percentage of simulation time during which agents are actively engaged in tasks rather than idle. Higher utilization indicates more efficient task allocation and less wasted time.
  • Task Allocation Efficiency: Calculates the number of tasks completed per distance traveled, indicating how effectively agents are navigating to and selecting tasks. Higher values suggest smarter path planning and better task assignment.
  • Error Rate: Measures the ratio of failed task attempts to total attempts. A lower error rate implies better trust calibration and decision-making, as agents avoid attempting tasks beyond their capabilities.
  • Throughput: The number of tasks completed per unit time (or per time step). Higher throughput indicates that tasks are being completed more rapidly.
  • Total Distance / Distance Per Task: The total distance traveled by all agents, as well as the average distance per completed task. These metrics capture navigation efficiency and resource expenditure.
  • Collision Rate: The frequency of collisions among agents. Fewer collisions generally indicate better coordination and route planning.
For each combination of trust model and scenario, we conducted 50 independent simulation runs to ensure statistical significance. Each run lasted for a maximum of 500 time steps or until all tasks were completed. We used the trust questionnaire from Yagoda and Gillian [34] as our theoretical framework for defining trust parameters, specifically focusing on measuring function and performance-related trust parameters, which include competence, ability, and dependability. The simulation logged key metrics at each step, including agent positions, trust levels, task completion times, and task success rates. This data allowed for a comparison of the trust models in terms of their impact on team performance and collaboration in a multi-agent SAR environment.
TABLE II: Performance metrics across trust models.
Trust Model Team Mean SR Std SR Mean CT Std CT Simple TS Med TS Comp TS Error AU AE
NoTrust 2H-2R 0.19 0.03 12.85 13.47 0.21 0.17 0.15 0.37 0.15 0.0868
NoTrust 5H-5R 0.41 0.07 27.42 19.47 0.39 0.46 0.39 0.54 0.15 0.0701
NoTrust 10H-10R 0.62 0.09 23.59 13.78 0.63 0.63 0.61 0.57 0.14 0.0530
Monir 2H-2R 0.19 0.03 16.03 18.49 0.19 0.19 0.18 0.39 0.14 0.0848
Monir 5H-5R 0.39 0.06 24.00 19.60 0.39 0.38 0.40 0.50 0.13 0.0647
Monir 10H-10R 0.60 0.09 23.34 15.03 0.58 0.61 0.60 0.55 0.12 0.0495
XuDudek 2H-2R 0.19 0.04 2.71 3.92 0.21 0.13 0.18 0.39 0.15 0.0822
XuDudek 5H-5R 0.41 0.07 4.92 9.91 0.43 0.36 0.42 0.48 0.15 0.0695
XuDudek 10H-10R 0.64 0.07 5.41 8.48 0.63 0.62 0.64 0.52 0.13 0.0535
GuoYang 2H-2R 0.20 0.04 15.02 20.94 0.20 0.20 0.25 0.47 0.15 0.0870
GuoYang 5H-5R 0.40 0.06 13.57 14.31 0.43 0.39 0.38 0.49 0.14 0.0677
GuoYang 10H-10R 0.62 0.07 18.66 13.33 0.69 0.61 0.59 0.58 0.12 0.0511
AliEtAl 2H-2R 0.19 0.04 7.57 13.29 0.18 0.20 0.24 0.38 0.15 0.0838
AliEtAl 5H-5R 0.40 0.06 11.83 12.19 0.39 0.38 0.41 0.50 0.15 0.0660
AliEtAl 10H-10R 0.62 0.08 11.54 9.71 0.57 0.65 0.62 0.55 0.14 0.0503
GuoEtAl 2H-2R 0.19 0.05 11.97 17.74 0.21 0.14 0.18 0.41 0.15 0.0857
GuoEtAl 5H-5R 0.41 0.08 14.69 13.22 0.41 0.41 0.41 0.49 0.14 0.0696
GuoEtAl 10H-10R 0.63 0.09 17.54 13.55 0.58 0.68 0.62 0.55 0.12 0.0530
ECT (Ours) 2H-2R 1.00 0.00 0.45 0.32 1.00 1.00 1.00 0.25 1.00 0.0321
ECT (Ours) 5H-5R 1.00 0.00 0.40 0.20 1.00 1.00 1.00 0.32 1.00 0.0191
ECT (Ours) 10H-10R 1.00 0.00 0.32 0.14 1.00 1.00 1.00 0.32 1.00 0.0125
Note: SR = Success Rate, CT = Completion Time, TS = Task Success (Simple, Medium, Complex), AU = Agent Utilization, AE = Allocation Efficiency.
@@ IV Results and Analysis
Table II summarizes the performance of seven trust models (NoTrust, Monir, XuDudek, GuoYang, AliEtAl, GuoEtAl, and ECT) across three team configurations (2H-2R, 5H-5R, 10H-10R). In the 2H-2R configuration, ECT achieves a perfect success rate of 1.00 and completes tasks in just 0.45 seconds on average, far exceeding the roughly 0.19–0.20 success rates and 7.57–16.03 second completion times of the other models. However, ECT’s impressive speed comes with substantially greater total travel distance (639.1 versus under 50 for most competitors), resulting in a notably lower allocation efficiency (0.0321). Its collision rate also stands at 0.834, higher than the 0.34–0.54 range observed in other models.
A similar pattern emerges with 5H-5R teams. ECT again posts a perfect success rate (1.00), compared to approximately 0.39–0.41 for the other models, and it maintains a remarkable average completion time of 0.40 seconds—an order of magnitude faster than the 11.83–27.42 seconds seen elsewhere. Yet ECT’s total distance climbs to 1079.6, far beyond the 100–125 range of the other approaches, while its allocation efficiency (0.0191) remains significantly lower, and its collision rate jumps to 9.363.
In the 10H-10R configuration, ECT continues to achieve a 1.00 success rate and completes tasks in only 0.32 seconds, whereas competing models display success rates in the 0.60–0.64 range and mean completion times ranging from 5.41 to 23.59 seconds. This level of performance again involves a total distance of 1664.2, substantially higher than the 242–251 range for the other models, and yields an allocation efficiency of 0.0125. ECT’s collision rate also rises to 42.76, far surpassing the next-highest rate of around 13.64, reflecting how its continuous high-level activity drives both perfect completion and intense agent interaction.
Overall, ECT’s trust-driven approach offers perfect success rates and extremely fast completion times for teams of all sizes. In scenarios where ensuring mission success and speed is paramount, such as time-critical search and rescue, these metrics may make ECT the optimal choice. However, the increases in total distance, collision rate, and reduced allocation efficiency highlight an inherent trade-off, as agents under ECT are more active but also more prone to collisions and resource inefficiencies. These results suggest that while ECT maximizes success and speed, future refinements could reduce collisions and excessive movement without compromising effectiveness.
V Conclusion
In this paper, we investigated the impact of trust on task allocation in multi-human multi-robot teams. We evaluated seven trust models, including N-Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, the Ali et al. Trust Model, the Guo et al. Trust Model, and our proposed ECT Model, across different team configurations (2H-2R, 5H-5R, and 10H-10R) in a simulated search and rescue scenario. Our findings indicate that the ECT model outperformed both the no-trust baseline and other trust models across most measured metrics. However, this performance came with a trade-off, leading to lower task allocation efficiency across all teaming formations. These results provide valuable insights into the role of trust in MH-MR task allocation and highlight the need for balancing efficiency with trust-based decision-making.